MATH 3014 Non-homogeneous DE (Undetermined Coefficient Method) 1. Solve the following differential equation. y - 4y + 3y =
Answer #1

If any confusion Please comment, I will help you, Thank you

Sop y - 44+3y= 2x Auxilary equation m2-4m + 3-0 m²-3m-m+3=0 mcom-3) -1cm-3)=0 (m-3)(m-1)=0 [m= 1,3] = Gesx c est + Gex and